When the bell in a clock tower rings with a sound of 473 Hz, a pigeon roosting in the belfry flies directly away from the bell. If the pigeon hears a frequency of 446 Hz, what is its speed?
Answer #1

f' = f (v - vo / v)

446 = 473 (343 - v0 / 343)

(446 / 473) * 343 = 343 - v0

323.4 = 343 - v0

v0 = 19.6 m/s